# Magnetic soliton molecules in binary condensates

### v1

@@ V Discussion
Future studies could investigate bound states in binary dipolar BECs, where long-range interactions and the presence of a roton-maxon spectrum may qualitatively affect the inter-soliton potentials and dynamics. Another interesting direction is the generalization to Rabi-coupled systems, which exhibit rich magnetic-soliton dynamics. Furthermore, one could study the crossover from the quasi-1D to the three-dimensional cigar-shaped regime. While magnetic solitons are expected to be more robust against three-dimensional instabilities than their single-component counterparts—since their size is governed by the spin-healing length, which grows large near the immiscibility threshold—it would be valuable to investigate the two-component analogue of snake instabilities and the formation of vortex rings. From an experimental perspective, oscillations of bound states, as well as collisions involving magnetic solitons and their molecular counterparts, could provide direct information about the interaction potentials, offering a route to test the predictions of this work.
@@ Appendix A Numerical calculation of single-soliton energy
$E_{\text{MS}}^{\text{GPE}}=\text{sign}(m_{\text{eff}})\left(E_{0}^{2L}/2-E_{0}\right),$ (14)
@@
Alternatively, recall that for like-sign solitons in the limit $\Delta x\to 0$, the inter-soliton potential $V$ [Eq. (11)] approaches the energy of a single magnetic soliton. In Fig.2(b), a single data point represents $V(\Delta x\to 0)$. This value cannot be obtained using realistic numerical grids. Instead, we employ a numerical workaround: during each step of imaginary-time evolution in a system of length $L$, the density is artificially constrained to zero for only a single grid point. This procedure yields excellent agreement with the single-soliton energy $E_{\text{MS}}^{\text{GPE}}$.

### v2

@@ V Discussion
Future studies could investigate bound states in binary dipolar BECs, where long-range interactions and the presence of a roton-maxon spectrum may qualitatively affect the inter-soliton potentials and dynamics. Another interesting direction is the generalization to Rabi-coupled systems, which exhibit rich magnetic-soliton dynamics. Furthermore, one could study the crossover from the quasi-1D to the three-dimensional cigar-shaped regime. While magnetic solitons are expected to be more robust against three-dimensional instabilities than their single-component counterparts—since their size is governed by the spin-healing length, which grows large near the immiscibility threshold—it would be valuable to investigate the two-component analog of snake instabilities and the formation of vortex rings. More generally, dark-antidark solitons also occur in imbalanced binary BECs, where the intraspecies interactions of the components differ [38]. We have confirmed that bound states also arise in such imbalanced systems (up to at least $g_{11}/g_{22}=1.2$; not shown); a systematic study of these bound states represents another interesting research direction. From an experimental perspective, oscillations of bound states, as well as collisions involving magnetic solitons and their molecular counterparts, could provide direct information about the interaction potentials, offering a route to test the predictions of this work.
@@ Appendix A Numerical calculation of single-soliton energy
$E_{\text{MS}}=\text{sign}(m_{\text{eff}})\left(E_{0}^{2L}/2-E_{0}\right),$ (14)
@@
Alternatively, recall that for like-sign solitons in the limit $\Delta x\to 0$, the inter-soliton potential $V$ [Eq. (11)] approaches the energy of a single magnetic soliton. In Fig.3(b), a single data point represents $V(\Delta x\to 0)$. This value cannot be obtained using realistic numerical grids. Instead, we employ a numerical workaround: during each step of imaginary-time evolution in a system of length $L$, the density is artificially constrained to zero for only a single grid point. This procedure yields excellent agreement with the single-soliton energy $E_{\text{MS}}$.